A researcher wanted to determine whether the average intelligence score for prisoners was different from the average intelligence score for the general population (μ = 100, σ = 15). Thirty-six prisoners on death row were administered an IQ test; the average IQ of the prisoners was 93. Use 0.05 significant level to test the null hypothesis.



1
Expert's answer
2022-05-08T14:01:35-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=100"

"H_a:\mu\not=100"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{93-100}{15/\sqrt{36}}=-2.8"

Since it is observed that "|z| = 2.8 > 1.96=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p =2P(Z<-2.8)= 0.00511," and since "p = 0.00511 < 0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 100, at the "\alpha = 0.05" significance level.